OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (6, 1), (5, 4)

OpenStudy (anonymous):

Hi, Yes! first find slope Finding equation of the line in standard form Ax + By = C Using the point-slope formula, m+=%28y%5B2%5D+-+y%5B1%5D%29%2F%28x%5B2%5D+-+x%5B1%5D%29 to find m (7, -3), (4, -8) m = -3-(-8) / 7-4 = 5/3 Using the standard slope-intercept form for an equation of a line y = mx + b where m is the slope and b the y-intercept. y = (5/3)x + b |using ordered pair Pt(7,-3) to solve for b -3 = (5/3)*7 + b - 3 - 35/3 = b -44/3 = b y = (5/3)x -44/3 |multiplying each term by 3 3y = 5x -44 5x - 3y = 44, equation of the line in standard form